This is a factoring question
A rectangular painting with a width of x centimeters has an area of x^2 + 50x square centimeters. Find a binomial that represents the length.

Answer by Nate(3500)

area = lw
area/w = l
(x^2 + 50x)/x = l
x(x + 50)/x = l
x + 50 = l